Suppose your company needs to raise $35.5 million and you want to issue 20-year bonds for this purpose. Assume the required return on your bond issue will be 8 percent, and you’re evaluating two issue alternatives: a 8.0 percent semiannual coupon bond and a zero coupon bond. Your company’s tax rate is 35 percent.

Requirement 1:
(a)

How many of the coupon bonds would you need to issue to raise the $35.5 million? (Do not round intermediate calculations. Enter the whole number for your answer, not millions (e.g., 1,234,567).)


  Number of coupon bonds   

(b)

How many of the zeroes would you need to issue? (Do not round intermediate calculations. Enter the whole number for your answer, not millions (e.g., 1,234,567). Round your answer to 2 decimal places (e.g., 32.16).)


  Number of zero coupon bonds   

Requirement 2:
(a)

In 20 years, what will your company’s repayment be if you issue the coupon bonds? (Do not round intermediate calculations. Enter your answer in dollars, not millions of dollars (e.g., 1,234,567).)


  Coupon bonds repayment $   

(b)

What if you issue the zeroes? (Do not round intermediate calculations. Enter your answer in dollars, not millions of dollars (e.g., 1,234,567). Round your answer to the nearest whole dollar amount (e.g., 32).)


  Zero coupon bonds repayment $   

Requirement 3:

Assume that the IRS amortization rules apply for the zero coupon bonds.

  

Calculate the firm’s aftertax cash outflows for the first year under the two different scenarios. (Do not round intermediate calculations. Input a cash outflow as a negative value and a cash inflow as a positive value. Enter your answers in dollars, not millions of dollars (e.g., 1,234,567). Round your answers to 2 decimal places (e.g., 32.16).)


  Coupon bond cash flow $   
  Zero coupon bond cash flow
Answer #1

1.a. Firstly, we need to calcuate the price of the coupon bond.

Yield (before tax) = 8% and after tax = 8*(1-0.35) = 5.2% = 0.052. Semiannual yield = 0.052/2.

Nper = 20*2 = 40 semiannual periods

Pmt = 8% of 1000 = 80 = 40 (semiannual)

FV = 1000

Price =pv(0.052/2,40,40,1000) = $1345.59

So. number of coupon bonds = 35,500,000/1345.59 = 26,382.41 bonds

1.b. Price of a zero coupon bond = FV/(1+yield)^n where yield after tax = 5.2% = 0.052/2 = 0.026, n= 40

So price of zero coupon bonds = 1000/1.026^40 = $358.18

Number of zero coupon bonds = 35,500,000/358.18 = 99,111.21 Bonds

2.a. In 20 years we will have to repay 40 coupon payments of $40 on each bond and the FV of 1000 at the end of 20 years.

Total bond repayment = 40*40*26,382.41 + 1000 * 26,382.41 = $68,954,266

2.b. For zero copon bonds. The total bond repayment will be 1000 * 99,111.21 = $99,111,210. There are no coupons in this case

Note: We have answered 4 sub parts of the question. Kindly repost to get the answers to remaining sub-parts